# Dynamics of Toxicity in Political Podcasts

### v1

@@ Abstract.
Toxicity in digital media poses significant challenges, yet little attention has been given to its dynamics within the rapidly growing medium of podcasts. This paper addresses this gap by analyzing political podcast data to study the emergence and propagation of toxicity, focusing on conversation chains—structured reply patterns within podcast transcripts. Leveraging state-of-the-art transcription models and advanced conversational analysis techniques, we systematically examine toxic discourse in over 30 popular political podcasts in the United States.
Our key contributions include: (1) creating a comprehensive dataset of transcribed and diarized political podcasts, identifying thousands of toxic instances using Google’s Perspective API, (2) uncovering concerning trends where a majority of episodes contain at least one toxic instance, (3) introducing toxic conversation chains and analyzing their structural and linguistic properties, revealing characteristics such as longer durations, repetitive patterns, figurative language, and emotional cues tied to anger and annoyance, (4) identifying demand-related words like ‘want,’ ‘like,’ and ‘know’ as precursors to toxicity, and (5) developing predictive models to anticipate toxicity shifts based on annotated change points.
Our findings provide critical insights into podcast toxicity and establish a foundation for future research on real-time monitoring and intervention mechanisms to foster healthier discourse in this influential medium.
Warning: Contains potentially abusive/toxic contents.
copyright: nonejournalyear: 2025conference: ; January; 2025isbn:
@@ 1. Introduction
Understanding and addressing toxicity in digital media is an ongoing challenge for researchers and practitioners alike. While much attention has been paid to platforms such as social media and online forums, comparatively less is known about the nature and dynamics of toxicity within the rapidly growing medium of podcasts. In this study, we aim to bridge this gap by collecting and analyzing political podcast data, focusing specifically on conversation chains – structured reply patterns within podcast transcripts – to study the emergence and propagation of toxicity in this medium.
@@
Podcasting has seen remarkable growth in recent years, emerging as a mainstream medium for information and entertainment. As of 2024, there are over 500 million podcast listeners worldwide, representing 23.5% of all internet users. In the United States alone, 47% of adults have listened to a podcast within the last month, a figure that has more than tripled over the past decade. Politics and government rank as the third most popular podcast topic, with 41% of listeners regularly tuning in to content in this category (Statista, 2024; Edison Research, 2024). This growth underscores the importance of podcasts as a platform for political discourse, but it also highlights their potential as vehicles for misinformation, conspiracies, and hate speech (Wirtschafter, 2023). Unlike traditional forms of media, podcasts remain largely unmoderated, further exacerbating the risks of unchecked toxicity. Understanding toxicity in podcasts is therefore not only academically significant but also crucial for ensuring healthy democratic discourse.
@@
Studying toxicity in podcasts presents unique challenges. Unlike text-based content, the audio format of podcasts makes them inherently difficult to monitor and analyze using traditional strategies. Transcribing large volumes of audio content incurs significant computational and financial costs, making large-scale studies impractical without significant resources. Further, audience interaction in podcasts is limited compared to social media platforms, where users can directly respond to or fact-check content. This lack of interactivity complicates efforts to understand how toxic narratives resonate with audiences. Moreover, analyzing toxicity within podcast transcripts requires not only identifying toxic language but also tracking how it ebbs and flows over time – examining what triggers toxic conversations and how they evolve within conversational structures. These complexities highlight why naïve approaches to studying podcast toxicity often fail to capture its multifaceted nature.
Existing research provides valuable starting points, but fails to address the problem in a comprehensive way. Advances in automatic transcription tools, such as Whisper, now allow for large-scale analysis of audio data. However, as previous work on datasets like RadioTalk (Beeferman et al., 2019) demonstrates, transcription errors and content ambiguity remain significant barriers, particularly for conversational and emotive formats such as podcasts. In addition, though the study of toxicity online is an active research area (Paz et al., 2020), there is little work on understanding conversational structures of toxicity and how it evolves (Zhang et al., 2024).
@@
In this paper, we introduce the concept of toxic conversation chains in podcasts. Using state-of-the-art transcription models and conversational analysis techniques, we explore how toxicity emerges and spreads in political podcasts. Our approach combines scalable transcription workflows with innovative analysis of conversational dynamics, providing novel insights into the structure and flow of toxic discourse in this medium.
Our work represents a critical step towards systematically analyzing toxicity in podcasts and lays the foundation for future research in this area.
Our key contributions and observations are as follows:
@@
  • We collect, transcribe and diarize a dataset of over 30 popular political podcasts from the US, reaching tens of millions of people, and identify thousands of instances of toxicity in these podcasts (as defined by Google’s Perspective API).
@@
  • We identify concerning trends of a majority of episodes from many popular podcasts containing at least one instance of toxicity.
@@
  • We build toxic conversation chains from the transcriptions where each chain has a highly toxic part which we call the anchor and a series of preceding and following parts to express the context of the anchor.
@@
  • Analysis of these message chains reveals that the anchor text is (a) longer in duration, (b) more repetitive, (c) incorporates figurative language such as metaphors and hyperboles, and (d) is closely tied to emotions like anger, fury, and annoyance.
  • Linguistic analysis reveals an intriguing observation: preceding words like ‘want,’ ‘like,’ and ‘know,’ which often express demands, can escalate into toxic conversations.
@@
  • We annotated the top 100 toxic conversation chains to identify change points where toxicity levels shift. Using these annotations, we developed a model to predict the temporal trajectory of these conversations. Our findings suggest that such automated methods hold significant potential for real-time monitoring and the development of effective intervention mechanisms in the future.
@@ 2. Related works
Toxicity on social media: Toxicity, broadly characterized as disrespectful, harmful, or unreasonable communication, has been a significant focus in social media research (Vogels, 2021). It manifests itself through harassment, hate speech, and polarized discourse, often perpetuated by bots and trolls that amplify its spread. Studies highlight that toxicity often combines with misinformation and polarization, forming a “cycle of online extremism” that exacerbates its societal impact (Vedeler et al., 2019).
Detection of toxicity: Defining and detecting toxicity requires nuanced approaches due to its context-sensitive nature. Techniques like BERT-based sentiment analysis have proven effective in detecting toxic content on platforms such as Twitter, leveraging large labeled datasets to improve classification accuracy (Mathew et al., 2022). Further, context-aware models are necessary to differentiate between benign negativity and harmful toxicity, particularly in multi-dimensional interactions (Gao and Huang, 2017). Perspective API (Lees et al., 2022) is a valuable tool for researchers to analyze toxicity of online content, allowing them to study the spread of hate speech, cyberbullying, and other harmful online behaviors in different languages and cultures.
Toxicity in podcasts: Unlike traditional social media, podcasts present unique challenges due to their audio form. Transcription is often necessary for textual analysis, but introduces errors and challenges such as speaker identification and role prediction (Clifton et al., 2020) that can skew results. Podcasts also allow for prolonged discussions, creating more opportunities for toxicity to escalate. Recent work has curated datasets to study toxicity in political podcasts, identifying patterns in conversational structures that promote or mitigate toxicity  (Litterer et al., 2024; Wirtschafter, 2021). Similarly, the authors in (Balci et al., 2024) analyze the political biases and content strategies of YouTube and Rumble podcasts, revealing a right-wing orientation in Rumble’s content and a more diverse perspective on YouTube. As a mitigation strategy, the authors in (Pathiyan Cherumanal et al., 2024) propose a novel approach to combat misinformation in podcasts using auditory alerts to notify listeners of potential misinformation in real-time.
Given this space, to the best of our knowledge, we are the first to look at toxicity on podcasts and to make use of the conversational aspects in podcasts. Our contributions are two fold: first is to show the prevalence of toxicity in popular political podcasts and the development of methods to look at toxicity in conversations rather than considering them as a one-off instance. Our methods and analysis extend beyond just podcasts and can be applicable in any conversational setting, which is prominent but highly under explored in NLP (Zhang et al., 2024).
@@ 3. Dataset
We construct the dataset based on channels list curated and shared by  (Wirtschafter, 2023). They identify popular political podcasts by analyzing Apple Podcasts’ top hundred lists from two specific time periods and selecting talk shows that discuss politics, policy, current events, or news. The process is expanded by including related political punditry podcasts from the ‘you might also like’ recommendations, creating a comprehensive sample for analysis. RSS feeds are used to download the wav files and the podcasts span around the years 2022-2023. The full list of 31 podcast shows we used in this paper are shown in Table 3 in the Appendix. We acknowledge that these channels might have a bias with over representation from right leaning sources since the original objective of the dataset curated by  (Wirtschafter, 2023) was to study conspiracies and misinformation. Having said that, care should be taken in interpreting the results particularly in the context of prevalence of toxicity, since our dataset only represents a small, biased subset of political podcasts.
@@
That said, even though the set of podcasts we consider is small, these shows might have disproportionate impacts and people who are in power might disproportionately listen to these. For e.g., a lot of January 6, 2021 US Capitol attack  [^1] https://www.britannica.com/event/January-6-U-S-Capitol-attack coordination and top down messages iterate through the podcasts of personalities like Steve Bannon. Hence, the claims made on these podcasts may have particularly consequential implications for broader public opinion and political discourse. So, as a part of this work, we try to understand the toxicity being driven through conversational chains so that such content can be auto-moderated and avoided. A detailed summary of the dataset over thirty-one podcast channels, and some metadata associated with them is shown in Appendix 3.
@@
We use Whisper (Radford et al., 2023) to transcribe and Pyannote [^2] https://pyannote.github.io/ to diarize and identify individual speakers (Bredin and Laurent, 2021). We filter out overlapping speech segments to avoid performance degradation in speaker clustering during diarization, which is accomplished using the same Pyannote model. The diarization process typically achieves high-quality results with over 90% accuracy, though some misclassifications do occur. For a detailed analysis of diarization quality, please refer to (Bredin and Laurent, 2021). Since our study focuses on conversation chains (see Section 4), it is crucial to maintain continuous speaker conversations without interruptions. In this context, the 90% accuracy is sufficient for our analysis, as we are examining conversation chains rather than individual speaker behaviors. When necessary, we outline the specific cleaning approaches used in subsequent sections.
@@
After processing, each episode in the dataset is made up of a number of speaker turn chunks, as returned by Pyannote library. Each chunk contains (a) start time (b) end time (c) unique speaker id, and, (d) diarized text. All these metadata are corresponding to these uneven chunks. As a next step, we further organize our data in the Section 4 for making conversational chains which are uniformly distributed to carry out our analysis.
@@ 4. Toxic conversation chains
In this section, we discuss the process employed to accurately calculate the toxicity scores and the subsequent formulation of toxic conversation chains using these toxicity scores.
4.1. Toxic score calculation
@@
To compute the toxicity of a speaker’s speech we pass the diarized text corresponding to the speech to Perspective API. [^3] https://developers.perspectiveapi.com/s/about-the-api-attributes-and-languages?language=en_US We use the benchmark Toxicity attribute, available with the API to obtain the toxicity scores. Since the Perspective API can handle a limited number of tokens we split the diarized conversation text corresponding to a speaker turn into chunks of 17 seconds since this is the median conversation time of a turn in our dataset (see Figure 1). We club 1-4 chunks (approximately one minute) into a segment. Each segment corresponds to a specific speaker identified by a speaker ID, a set of (at most) four chunks, the start and end timestamps and the diarized text corresponding to the chunks in the segment. The toxicity score of a segment is set to the maximum of the toxicity score across these chunks within the segment. If a speaker turn exceeds the duration limit of a segment, it is broken into the required number of contiguous segments. We perform all our analysis at the level of these one minute segments so that the observations are statistically meaningful and not over dominated by a speaker speaking over a long stretch.
Figure 2 shows the top 10 podcasts with the highest percentage of episodes containing at least one instance of toxic conversations. Certain shows such as Get Off My Lawn Podcast w/ Gavin Mcinnes, The New Abnormal, and Louder with Crowder have at least one instance of a toxic conversation in all of their episodes. Table 3 shows the full statistics for all the shows in our dataset. It is really surprising to see instances of toxicity in hundreds of episodes even on shows with tens of millions of listeners (e.g. The Dan Bongino Show).
4.2. Toxic chain formulation
@@
A toxic chain constitutes an anchor segment, the preceding ten and the following ten segments of the anchor segment (see Figure 1). A segment is identified as an anchor segment if its toxicity score is 0.7 or higher. This threshold is set as per the guidelines noted in documentation of the Perspective API. [^4] https://support.perspectiveapi.com/s/about-the-api-score?language=en_US We take the previous and the next ten segments to understand what leads to the rise in toxicity and what happens after a peak toxic threshold has been reached. Anchoring on this threshold, we perform our analysis on a total of 8,634 chains. A set of representative examples of toxic conversation chains are provided in Figure 8 & Appendix 11.
@@
Figure 2. Top 10 podcast shows with most amount of toxic content. The bars show the percentage of episodes containing at least one toxic conversation.
4.3. Distribution of the toxic chains across podcasts
We compute the distribution of the toxic chains across the different podcast channels. Figure 3 shows the distribution of top ten channels. The top five contributing podcast channels are Louder with Crowder, Mark Levin Podcast, The New Abnormal, The Dan Bongino Show and Get Off My Lawn Podcast w/ Gavin McInnes in decreasing order of contribution aggregating to a total of 63.6%. Together, these podcasts potentially reach tens of millions of listeners every week.555e.g. Just Dan Bongino and Mark Levin podcasts are extremely popular with over 20 million weekly listeners. See https://bit.ly/3ZEArcS Many of these channels have been reported to be hosted and attended by people making homophobic, and racist remarks. [^6] https://tinyurl.com/louder-crowder, [^7] https://tinyurl.com/dan-bingo-show1, [^8] https://tinyurl.com/dan-bingo-show2
@@ 4.3. Distribution of the toxic chains across podcasts
Figure 3. Distribution of toxic conversation chains across the podcast channels. Percentage contribution for top 10 podcast channels are shown.
5. Analysis of the toxic conversation chains

### v2

# Toxicity Begets Toxicity: Unraveling Conversational Chains in Political Podcasts

@@ Abstract.
Tackling toxic behavior in digital communication continues to be a pressing concern for both academics and industry professionals. While significant research has explored toxicity on platforms like social networks and discussion boards, podcasts—despite their rapid rise in popularity—remain relatively understudied in this context. This work seeks to fill that gap by curating a dataset of political podcast transcripts and analyzing them with a focus on conversational structure. Specifically, we investigate how toxicity surfaces and intensifies through sequences of replies within these dialogues, shedding light on the organic patterns by which harmful language can escalate across conversational turns. Warning: Contains potentially abusive/toxic contents. [^1] Accepted as regular paper (main track) at ACM Multimedia 2025
Toxic conversation chains, podcasts, transcripts, change point detection, toxicity begets toxicity
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754553isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Speech recognitionccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Language resourcesccs: Computing methodologies Temporal reasoningccs: Computing methodologies Information extractionccs: Social and professional topics Hate speechccs: Social and professional topics Political speech
@@ 1. Introduction
Understanding and addressing toxicity in digital media is an ongoing challenge for researchers and practitioners alike. While much attention has been paid to platforms such as social media and online forums, comparatively less is known about the nature and dynamics of toxicity within the rapidly growing medium of podcasts. In this study, we aim to bridge this gap by collecting and analyzing political podcast data, focusing specifically on conversation chains – structured reply patterns within podcast transcripts – to study the emergence and propagation of toxicity in this medium, naturally unfolding how toxicity begets toxicity through conversation chains.
@@
Rise in the popularity of podcasts: Podcasting has seen remarkable growth in recent years, emerging as a mainstream medium for information and entertainment. As of 2024, there are over 500 million podcast listeners worldwide, representing 23.5% of all internet users. In the United States alone, 47% of adults have listened to a podcast within the last month, a figure that has more than tripled over the past decade. Politics and government rank as the third most popular podcast topic, with 41% of listeners regularly tuning in to content in this category [^2] https://www.statista.com/statistics/270365/audio-podcast-consumption-in-the-us/
https://www.edisonresearch.com/the-podcast-consumer-2024-by-edison-research/
. This growth underscores the importance of podcasts as a platform for political discourse, but it also highlights their potential as vehicles for misinformation, conspiracies, and hate speech (Wirtschafter, 2023). Unlike traditional forms of media, podcasts remain largely un-moderated, further exacerbating the risks of unchecked toxicity. Understanding toxicity in podcasts is therefore not only academically significant but also crucial for ensuring healthy democratic discourse.
@@
Need of the hour: A recent post by Meta [^3] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes/ highlights the difficulty of real-time proactive content moderation and one such massive outrage that recently broke out in India [^4] https://www.ndtv.com/feature/indias-got-latent-row-who-said-what-as-samay-raina-forced-to-remove-all-youtube-episodes-7708513 is a prime example of the need of the hour. Studying toxicity in podcasts presents unique challenges thus limiting the effectiveness of real-time proactive content moderation. Unlike one-off contents like text-based posts, memes, and GIFs, the audio format of podcasts makes them inherently difficult to monitor and analyze using traditional strategies, primarily due to their continuous nature. Transcribing large volumes of audio content incurs significant computational and financial costs, making large-scale studies impractical without significant resources. Further, audience interaction in podcasts is limited compared to social media platforms, where users can directly respond to or fact-check content. This lack of interactivity complicates efforts to understand how toxic narratives resonate with audiences. Moreover, analyzing toxicity within podcast transcripts requires not only identifying toxic language but also tracking how it ebbs and flows over time – examining what triggers toxic conversations and how they evolve within conversational structures. These complexities highlight why naïve approaches to understand toxicity in podcasts often fail to capture its multifaceted nature. Existing research provides valuable starting points, but fails to address the problem in a comprehensive way. Advances in automatic transcription tools, such as Whisper, now allow for large-scale analysis of audio data. However, as previous work on datasets like RadioTalk (Beeferman et al., 2019) demonstrates, transcription errors and content ambiguity remain significant barriers, particularly for conversational and emotive formats such as podcasts. In addition, though the study of toxicity online is an active research area (Paz et al., 2020), there is little work on understanding conversational structures of toxicity and how it evolves (Zhang et al., 2024).
@@
Contributions: In this paper, we introduce a fresh research statement for audio content moderation proposing a novel concept of toxic conversation chains and utilizing them in the study of how toxicity begets toxicity. This formulation allows us to understand the growth, spread and decay of toxicity in conversations. Using state-of-the-art transcription models and conversational analysis techniques, we explore how toxicity emerges, spreads and decays in political podcasts. Our approach combines scalable transcription workflows with innovative analysis of conversational dynamics, providing crucial insights into the structure and flow of toxic discourse in this medium. Our work represents a critical step towards systematically analyzing toxicity in podcasts and lays the foundation for future research in this area. Our key contributions and observations are as follows:
@@
(i) Collecting, transcribing and diarizing a dataset over 52 popular political podcasts (31 right- and 21 left-leaning) from the US, reaching tens of millions of people, and identify thousands of instances of toxicity in these podcasts. The dataset encompasses a total of 12,322 episodes from these channels (9,166 from right- and 3,156 from left-leaning podcasts).
@@
(ii) Identifying concerning trends of a majority of episodes from many popular podcasts containing at least one instance of toxicity.
@@
(iii) Introducing toxic conversation chains, built from the transcriptions where each chain has a highly toxic part which we call the anchor and a series of preceding and following parts to express the context of the anchor.
@@
(iv) Analysis of these message chains reveals that the anchor text is (a) longer in duration, (b) more repetitive, (c) incorporates figurative languages such as metaphors and hyperboles, and (d) is closely tied to emotions like anger, fury, and annoyance.
@@
(v) Annotating a novel dataset based on the premise of toxicity begetting toxicity from top 200 toxic conversation chains (100 each from right- and left-leaning channels) to identify change points where toxicity levels shift. Using these annotations, we proposed a setup to predict the temporal trajectory of these conversations. Our findings suggest that such automated methods hold significant potential for real-time monitoring and the development of effective intervention mechanisms in the future [^5] Supplementary material, dataset and code – https://github.com/hate-alert/ToxicityBegetsToxicity-Audio..
@@ 2. Related works
Detection of toxicity: Defining and detecting toxicity requires nuanced approaches due to its context-sensitive nature. Techniques like BERT-based sentiment analysis have proven effective in detecting toxic content on platforms such as Twitter, leveraging large labeled datasets to improve classification accuracy (Mathew et al., 2022). Further, context-aware models are necessary to differentiate between benign negativity and harmful toxicity, particularly in multi-dimensional interactions (Gao and Huang, 2017) and across different modalities (Rizwan et al., 2025a). Perspective API (Lees et al., 2022) is a valuable tool for researchers to analyze toxicity of online content, allowing them to study the spread of hate speech, cyberbullying, and other harmful content in different languages.
Toxicity in podcasts: Unlike traditional social media, podcasts present unique challenges due to their audio form. Transcription is often necessary for textual analysis, but introduces errors and challenges such as speaker identification and role prediction (Clifton et al., 2020) that can skew results. Podcasts also allow for prolonged discussions, creating more opportunities for toxicity to escalate. Recent work has curated datasets to study toxicity in political podcasts, identifying patterns in conversational structures that promote or mitigate toxicity  (Litterer et al., 2024; Wirtschafter, 2021). Similarly, the authors in (Balci et al., 2024) analyze the political biases and content strategies of YouTube and Rumble podcasts. As a mitigation strategy, the authors in (Pathiyan Cherumanal et al., 2024) propose an approach to combat misinformation in podcasts using auditory alerts to notify listeners of potential misinformation in real-time.
LLMs and audio-LLMs: In recent years, the advent of transformers (Vaswani et al., 2023) has led to the development of various state-of-the-art LLMs (Minaee et al., 2024) which have proven their capabilities in very complex applications. However, research on audio-LLMs for conditional generation given an audio input is limited (Chu et al., 2024; OpenAI, 2024) [^6] https://platform.openai.com/docs/guides/audio. Recently researchers have explored LLMs for change point detection as well (Dong et al., 2024). However, to the best of our knowledge, audio-LLMs have not been employed for such a complex task of detecting change points in a toxic conversation before.
Present work: To the best of our knowledge, we are the first to look at toxicity in podcasts and use the conversational aspects in podcasts. Our contributions are two fold: first is to show the prevalence of toxicity in popular political podcasts and the development of methods to look at toxicity in conversations rather than considering them as a one-off instance.
@@ 3. Dataset
Curation: We construct the dataset based on the channels list curated and shared by  (Wirtschafter, 2023) and also from the following collection hosted at https://politicalpodcastproject.shinyapps.io/dataset/. They identify popular political podcasts by analyzing and selecting talk shows that discuss politics, policy, current events, or news. We further expand the channels by including related political punditry podcasts from the ‘you might also like’ recommendations, creating a comprehensive sample for analysis. RSS feeds and the link to episodes from the above collection are used to download the audio files, and the podcasts span over the years 2022-2023. The full list of 52 podcast channels (31 right- and 21 left-leaning), making a total of 12,322 episodes (with 9,166 from right- and 3,156 from left-leaning channels) is noted in the Appendix [^7] Care should be taken in interpreting the results particularly in the context of prevalence of toxicity, since our dataset only represents a small subset of political podcasts.. Even though the set of podcasts we consider is small, these shows might have disproportionate impacts and people who are in power might disproportionately listen to these. For e.g., a lot of January 6, 2021 US Capitol attack [^8] https://www.britannica.com/event/January-6-U-S-Capitol-attack coordination and top down messages iterate through the podcasts of personalities like Steve Bannon. Hence, the claims made on these podcasts may have particularly consequential implications for broader public opinion and political discourse. So, as a part of this work, we try to understand the toxicity being driven through conversational chains so that such content can be auto-moderated and avoided.
@@
Processing: We use Whisper (Radford et al., 2023) to transcribe and Pyannote [^9] https://pyannote.github.io/ to diarize and identify individual speakers (Bredin and Laurent, 2021). We filter out overlapping speech segments to avoid performance degradation in speaker clustering during diarization, which is accomplished using the same Pyannote model. The diarization process typically achieves high-quality results with over 90% accuracy, though some misclassifications do occur. For a detailed analysis of diarization quality, please refer to (Bredin and Laurent, 2021). Since our study focuses on conversation chains (see Section 4), it is crucial to maintain continuous speaker conversations without interruptions. In this context, the 90% accuracy is sufficient for our analysis, as we are examining conversation chains rather than individual speaker behaviors. When necessary, we outline the specific cleaning approaches used in subsequent sections.
@@
Metadata: After processing, each episode in the dataset is made up of a number of speaker turn chunks, as returned by Pyannote library. Each chunk contains (a) start time (b) end time (c) unique speaker id, and (d) diarized text. All these metadata correspond to these uneven chunks. As a next step, we further organize our data in the Section 4 for making conversational chains that are uniformly distributed to carry out our analysis. A detailed summary of the dataset over 52 podcast channels, and some more metadata associated with them, are presented in the Appendix.
@@ 4. Toxic conversation chains
In this section, we discuss the process employed to accurately calculate the toxicity scores and the subsequent formulation of toxic conversation chains using these toxicity scores. Note that our work is centered to understand and intervene upon the trends of toxicity across audio podcasts. Hence, we use state-of-the-art toxicity scores generator as an aid to study conversational trends.
@@
Segments: To compute the toxicity of a speaker’s speech we pass the diarized text corresponding to the speech to Perspective API. [^10] https://developers.perspectiveapi.com/s/about-the-api-attributes-and-languages We use the benchmark Toxicity attribute, available with the API to obtain the toxicity scores. Since the Perspective API can handle a limited number of tokens we split the diarized text corresponding to a speaker turn into chunks of 17 seconds since this is the median conversation time of a turn in our dataset (see Figure 1). We club 1-4 chunks (approximately one minute) into a segment. Each segment corresponds to a specific speaker identified by a speaker ID, a set of (at most) four chunks, the start and end timestamps and the diarized text corresponding to the chunks in the segment. The toxicity score of a segment is set to the maximum of the toxicity score across these chunks within the segment. If a speaker turn exceeds the duration limit of a segment, it is broken into the required number of contiguous segments. All our analyses are at the level of these one minute segments so that the observations are statistically meaningful and not over dominated by a speaker speaking over a long stretch.
@@
Conversation chain: A toxic chain constitutes an anchor segment, the preceding ten and the following ten segments of the anchor segment (see Figure 1). A segment is identified as an anchor segment if its toxicity score is 0.7 or higher. This threshold is set as per the guidelines noted in documentation of the Perspective API [^11] https://developers.perspectiveapi.com/s/about-the-api-score (refer to ‘Supporting human moderators’ subsection). We take the previous and the next ten segments to understand what leads to the rise in toxicity and what happens after a peak toxic threshold has been reached. We manually tune to ten previous and next segments, as this gives a perfect amount of back-and-forth context around the anchor point. Choosing fewer segments leads to the loss of context and incorporating more segments was not beneficial. Further, this selection is just precise for the manual CPD annotation as well (refer to Section 6). Anchoring on this threshold, we perform our analysis on a total of 8,634 chains from right- and 7,124 chains from left-leaning channels. A set of representative examples of toxic conversation chains is provided in Figure 4 & in the Appendix. Further, extended analysis over the properties of these chains (especially for the left-leaning ones) is presented in the Appendix.
Robustness of Perspective API scores: To establish the robustness of toxicity scores assigned by Perspective API, we perform additional validation. Specifically, we use two different types of models and further verify the anchor segment of toxic conversation chains via human annotation. We perform these additional robustness experiments on the right-leaning toxic conversation chains (results have exactly similar trends for the left-leaning channels and hence are not shown to avoid repetition). Following are the key results:
(i) Detoxify library generated scores: We run two different Toxic-BERT models (original & unbiased) [^12] https://dataloop.ai/library/model/unitary_toxic-bert/ and utilize their Toxicity attribute scores. Both of these models align with nearly 93% of the cases where Perspective API assigns toxicity scores higher than 0.7.
(ii) Classification using GPT-4o: We prompt GPT-4o to classify whether the diarized text of the anchor segment is toxic or not and provide the definition for toxicity from the Perspective API manual as additional context in the prompt. GPT-4o classified 86% of the total anchor segments to be toxic aligning with the cases where the toxicity score from Perspective API is greater than 0.7. Please refer to the Appendix for the detailed prompt.
(iii) Human evaluation: We conduct binary annotation task (toxic or non-toxic) on anchor segment’s diarized text over 100 randomly selected chains using two expert annotators and found 86 of them to be either toxic, vulgar or attacking in nature to be spoken on a live-streaming show. The results are obtained with a high inter-annotator agreement of 0.69 based on Cohen’s $\kappa$. During annotation, annotators were asked to stick to the definition of Toxicity attribute as per the Perspective API documentation.
Based on the obtained results, we henceforth conclude that a Perspective API score of 0.7 is a reasonable threshold to flag the anchor text of a chain to be toxic.
@@
Figure 2. Top 10 podcast shows with most amount of toxic content for each leaning, i.e. right and left. The bars show the percentage of episodes containing at least one toxic conversation.
@@
Figure 3. Distribution of toxic conversation chains across the podcast channels for each leaning, i.e. right and left. Percentage contribution for Top 10 podcast channels are shown.
Figure 4. Example: Since it is not feasible to illustrate all segments, one among the previous and next segments is shown along with the anchor segment from the toxic conversation chain. Toxic texts in the anchor segment are marked in red. Note: start and end times are in seconds.
Most toxic channels: Figure 2 shows the top 10 podcasts with the highest percentage of episodes containing at least one instance of toxic conversations for both right- and left-leaning channels. Certain shows such as Get Off My Lawn Podcast w/ Gavin Mcinnes, The New Abnormal, and Louder with Crowder from right-leaning and Lovett or Leave It, Mea Culpa, Krystal Kyle and Friends, Hysteria and Lady Dont Take No from left-leaning channels, have at least one instance of a toxic conversation in almost all of their episodes. Tables in the Appendix show the full statistics for all the shows in our dataset. It is quite surprising to see instances of toxicity in hundreds of episodes, even on shows with tens of millions of listeners (e.g. The Dan Bongino Show).
Distribution of the toxic chains across podcasts: We compute the distribution of the toxic chains across the different podcast channels. Figure 3 shows the distribution of top ten channels for both right- and left-leaning channels. The top five contributing right-leaning podcast channels are Louder with Crowder, Mark Levin Podcast, The New Abnormal, The Dan Bongino Show and Get Off My Lawn Podcast w/ Gavin McInnes in decreasing order of contribution, aggregating to a total of 63.6%. From the left-leaning channels, Mea Culpa and Lovett or Leave it contribute nearly 45.4% to toxic conversation chains. Together, these podcasts potentially reach tens of millions of listeners every week131313e.g., Just Dan Bongino and Mark Levin podcasts are extremely popular with over 20 million weekly listeners (see https://bit.ly/3ZEArcS). Many of these channels have been reported to be hosted and attended by people making homophobic, and racist remarks [^14] https://tinyurl.com/louder-crowder, [^15] https://tinyurl.com/dan-bingo-show1, [^16] https://tinyurl.com/dan-bingo-show2.
5. Analysis of toxic conversation chains